### v3

# SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization

Abstract
Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited. To further enhance the efficiency of attention computation compared to SageAttention while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques. First, we propose to quantize matrixes $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$. Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$. The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 5x on RTX4090, respectively. Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for large language processing, image generation, and video generation. The codes are available at https://github.com/thu-ml/SageAttention.
footnotetext: * Equal Contribution.
Figure 1: The upper left figure shows the kernel speedup on RTX4090 GPU. The upper right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the needle-in-a-haystack task (gkamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU. The figure below shows two videos generated by CogvideoX (1.5-5B) using FlashAttention2 and SageAttention2 on RTX4090, showing that SageAttention2 accelerates generation 1.8x with no video quality loss.
1 Introduction
Due to the quadratic time complexity of attention, efficiently implementing it becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024). Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., ; Xiao et al., 2023b, 2024; Chen et al., 2023; Jiang et al., 2024; Venkataramanan et al., 2023; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks. Widely used attention methods optimize attention by exploiting hardware capacities to enhance the computation speed, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2023), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2024). Those works do not omit the computation of attention over parts of the context and achieve impressive speed and precision performance across various models and tasks.
Motivation. For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $\widetilde{P}V$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $\widetilde{P}V$. Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers. SageAttention achieves 2 $\times$ and 2.7 $\times$ speedup than FlashAttention2 and xformers and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models. However, SageAttention has two weaknesses. (W1) INT8 Matmul achieves only half the speed of INT4. (W2) FP16 Matmul with FP16 accumulators is only compatible with RTX4090 and RTX3090 GPUs. To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose SageAttention2, which quantizes $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
Challenges. Quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8 presents significant challenges. For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video (See Figure 2), and Llama2 only achieves a random-guessing-level accuracy of 25% on MMLU. After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values. (C2) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. This will lead to an accuracy loss of $\widetilde{P}V$.
Figure 2: An example of quantizing Q, K to INT4 from CogvideoX.
Our approach. To address (C1), we first discovered that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization. Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead. Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU thread and memory layout of matrices as dictated by the PTX mma instructions. This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale. This approach achieves much better accuracy performance than per-block quantization with no additional dequantization overhead. Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and propose an effective method to remove these outliers in $Q$. Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$. Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation. To address (C2)—the accuracy loss from using a 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$—, we propose to use an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$. This method confines the errors to the block range. Additionally, we design an optional technique to enhance the accuracy of the 22-bit accumulator. Specifically, we could smooth $V$ by subtracting the average value of its channel dimension and adding the subtracted item to the attention output to maintain the correctness.
Performance. Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs. This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 5x, respectively. Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture, i.e., H100 and H800. We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2. Across all models and tasks, SageAttention2 can be directly adopted in a plug-and-play manner with negligible loss in model performance.
2 Preliminary
2.1 FlashAttention
The computation of self-attention can be formulated as follows: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. The matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$, $P$ are $N\times N$. While $d$ is typically small, e.g., 64 or 128, $N$ can be millions. Therefore, the $N\times N$ matrices $(S,P)$ are much larger than $(Q,K,V)$, and a naive implementation suffers from the huge amount of global memory IO for $(S,P)$ reads/writes. FlashAttention (Dao, 2023) proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block size of $b_{q}$, $b_{k}$, $b_{v}$, respectively, where $b_{k}$ = $b_{v}$. Then, to avoid the memory IO for $(S,P)$, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
First, for each block of $\{K_{i}\},\{V_{i}\}$, it computes the following equations iteratively:
$S^{j}_{i}=Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m^{j}_{i},\widetilde{P}^{j}_{i})=\tilde{\sigma}(m^{j-1}_{i},S^{j}_{i}),~{}~{}l_{i}^{j}=$ $\exp(m^{j-1}_{i}-m^{j}_{i})l_{i}^{j-1}+\mathrm{rowsum}(\widetilde{P}^{j}_{i}),$
$O^{j}_{i}=\mathrm{diag}\left(\exp(m^{j-1}_{i}-m^{j}_{i})\right)O^{j-1}_{i}$ $+\widetilde{P}^{j}_{i}V_{j}$
Where $m_{i}^{j}$ and $l_{i}^{j}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively. The $\tilde{\sigma}()$ is an online softmax operator: $m^{j}_{i}=\max\{m^{j-1}_{i},\mathrm{rowmax}(S^{j}_{i})\},~{}\widetilde{P}_{j}^{i}=\exp(S^{j}_{i}-m^{j}_{i})$.
Finally, the output $O_{i}$ can be computed by $O_{i}=\mathrm{diag}(l_{i}^{j})^{-1}O_{i}^{j}$.
2.2 Quantization
A matrix multiplication $C=AB$ can be accelerated with quantization as:
$(\delta_{A},\hat{A})=\psi(A),~{}~{}(\delta_{B},\hat{B})=\psi(B),~{}~{}\hat{C}=\hat{A}\hat{B},$ $~{}~{}C=\psi^{-1}_{\delta_{A}\delta_{B}}(\hat{C})$ (1)
$\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision. We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$. The actual matrix multiplication $\hat{A}\hat{B}$ is carried in low precision. In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones. Many quantizers depend on the numerical format and granularity, e.g., how many elements share a common scale factor. For example, an INT4 per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$. Likewise, per-token quantizer assigns a scale factor for each token of a tensor: $\hat{A}[i,:]=\lceil A[i,:]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|{A[i,:]}|)/7$. Also, per-channel quantizer assigns a scale factor for each channel of the tensor, i.e., along the channel dimension: $A[:,i]=\lceil A[:,i]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|A[:,i]|)/{7}$. The dequantization process is an element-wise scaling. For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}*\delta_{A}*\delta_{B}$.
2.3 SageAttention
Based on the block tiling in FlashAttention2 (Dao, 2023), SageAttention (Zhang et al., 2024) quantize $Q,K$ to INT8 in a per-block granularity. Moreover, to keep the quantization accuracy, SageAttention proposes to smooth $K$ first:
$K=K$ $-\text{mean}(K)$
$\hat{Q}_{i}=\lceil Q_{i}/\delta_{Q}\rfloor,~{}~{}\delta_{Q}=\max(|{Q_{i}|})/127,$ $~{}~{}\hat{K}_{j}=\lceil K_{j}/\delta_{K}\rfloor,~{}~{}\delta_{K}=\max(|{K_{j}|})/127$
$S_{ij}=$ $Q_{i}K_{j}^{\top}*\delta_{Q}*\delta_{K}$
Where $Q_{i},K_{j}$ are the tilled block in FlashAttention, and $\text{mean}(K)$ is the average value of the channel dimension of $K$. Additionally, SageAttention keeps $\widetilde{P}V$ as FP16 and uses FP16 Matmul with an FP16 accumulator for $\widetilde{P}V$. However, FP16 Matmul with an FP16 accumulator only has a speedup effect on RTX4090 and RTX3090 GPUs.
3 SageAttention-2
Figure 3: Workflow of SageAttention2. 1 Smooth Q,K,V. 2 A GEMV to obtain $\Delta S$. 3 Per-thread quantize Q,K and per-channel quantize V. 4 Perform the SageAttention2 kernel. 5 Correct the output.
3.1 Formulation
Based on the introduction of FlashAttention and quantization presented in Section 2, we now describe the quantized attention approach we developed.
Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q},~{}~{}\overrightarrow{Q}_{m})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{Q}}}(Q),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{K}}}(K),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V},\overrightarrow{V}_{m})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{V}}}(V)$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\Delta S=\overrightarrow{Q_{m}}K^{\top}}}$ (2)
Attention: $S={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi^{-1}_{\delta_{Q}\delta_{K}}(\hat{Q}\hat{K}^{\top})+\Delta S}},~{}~{}(m^{\prime},\widetilde{P})=\tilde{\sigma}(m,S),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{P},\hat{P})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{P}}}(\widetilde{P})$
$O=\mathrm{diag}\left(\exp(m^{\prime}-m)\right)O+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi^{-1}_{\delta_{P}\delta_{V}}(\hat{P}\hat{V})}}$ (3)
$\phi_{Q}$, $\phi_{K}$, $\phi_{V}$ are three transformations to obtain quantized $Q$, $K$, and $V$, which we will discuss in subsequent sections. For simplicity, we omit all superscripts and subscripts, but the matrices used in attention are still tiles, and the computation is still organized as FlashAttention described in Section 2.1. Compared to the original full-precision version, as shown in Eq. 23, SageAttention2 adds quantizers to $Q,K,P,V$ and dequantizers to the product to accelerate both Matmuls of $QK$ and $\widetilde{P}V$.
Figure 4: Typical examples of tensors’ data distribution in attention.
Table 1: End-to-end metrics comparison of different quantization methods, where $Q,K$ are quantized into INT4, while $\widetilde{P},V$ stay in full precision.
Q, K Smoothing (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikeText) $\downarrow$ CogVideoX (vqa-a) $\uparrow$ CogVideoX (vqa-t) $\uparrow$
Full-Precision - 81.5% 6.013 77.605 75.360
INT4 Quantization 72.6% 11.698 27.114 24.670
80.8% 6.219 77.276 75.147
Figure 5: An example of quantized value distribution of $Q$ before and after smoothing $Q$.
3.2 Smooth Q
The representative range for quantization of INT4 is notably restrictive, i.e., only $152^{4}-1=$ 15 values. This limitation significantly degrades performance when quantizing $Q,K$ into INT4 in attention. For example, using INT4 to quantize $Q,K$ can cause the perplexity of Llama3.1 on WikiText to increase by more than 90%, and the quality of videos generated by CogvideoX decrease about 3x (See Table 1). We analyze the data distribution of $Q,K$ in real models. For example, we find that Llama3.1 and CogvideoX show distinct channel-wised outliers in $Q,K$ (See Figure 4). While per-channel quantization could mitigate the quantization error caused by such outliers, this method is impracticable for $Q,K$ because quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$.
To address this challenge, we propose a method to eliminate the quantization error brought by the outliers in $Q,K$:
$\overrightarrow{Q_{m}}=\mathrm{mean}(Q),~{}~{}\gamma(Q)=$ $Q-\overrightarrow{Q_{m}},~{}~{}\gamma(K)=K-\mathrm{mean}(K),$
$S=QK^{\top}=$ $\gamma(Q)\gamma(K)^{\top}+\overrightarrow{Q_{m}}\gamma(K)^{\top}$
$\phi_{Q}(Q)=\psi_{Q}\circ\gamma(Q),~{}\overrightarrow{Q_{m}}$ $,~{}~{}~{}~{}~{}\phi_{K}(K)=\psi_{K}\circ\gamma(K)$ (4)
Where $\mathrm{mean}(K)$ and $\gamma(K)$ have been discussed in SageAttention (Zhang et al., 2024), and the transformation for $K$ does not change the attention score $\widetilde{P}$. $\mathrm{mean}(Q)=\frac{1}{N}\sum_{t=1}^{N}Q[t,:]$ is a vector with a shape $1\times d$. For the transformation of $Q$, we will perform a GEMV (general matrix-vector multiplication) between $\overrightarrow{Q_{m}}$ and $K^{\top}$, i.e., $\overrightarrow{Q_{m}}K^{\top}$ as $\Delta S$. This $\Delta S$ will be added to $S$ during the attention computation. Finally, the transformation from full-precision $Q,K$ to quantized $\hat{Q},\hat{K}$ can be expressed as shown in Eqution 4, where $\psi_{Q}$, $\psi_{K}$ is two quantizers for $Q$ and $K$. In other words, full-precision $Q,K$ are substracted with the means of the channel dimension before being quantized.
Figure 5 shows an example from CogvideoX of the distribution of quantized $Q$ with and without smoothing Q. We can find that with smoothing Q, the range of INT4 is utilized more uniformly and fully. Table 1 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 (Dubey et al., 2024) and CogvideoX (Yang et al., 2024). The results demonstrate that smoothing Q+K offers significant accuracy benefits. Also, Table 9 and Table 9 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.
Figure 6: An example of per-thread quantization. The left figure shows the correspondence between the quantization scales and the tokens in each GPU warp. The right figure shows the correspondence between quantization tokens and GPU threads in a MMA.m16n8k64 instruction, showing that each GPU thread only corresponds to one quantization scale in $\delta_{Q}$ and $\delta_{K}$ in dequantization.
Table 2: Average accuracy across all layers of CogvideoX using different quantization granularities.
Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
Per-token 99.45% 0.0649 0.0335
Per-thread 99.45% 0.0622 0.0313
Per-block 98.03% 0.1492 0.0744
Per-tensor 97.15% 0.1800 0.0865
Table 3: Worst accuracy across all layers of CogvideoX using different quantization granularities.
Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
Per-token 96.76% 0.1916 0.0775
Per-thread 96.72% 0.1932 0.0776
Per-block 90.68% 0.3615 0.1490
Per-tensor 85.85% 0.4687 0.2261
3.3 INT4 Per-thread Quantization
SageAttention uses per-block quantization, which quantizes each block of $Q$ and $K$ for a GPU streaming processor. Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the overhead of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$. However, quantizing $Q$ and $K$ to INT4 demands a more accurate quantization granularity. We propose per-thread quantization, a more precise and granular quantization approach than the per-block quantizer, also without the additional overhead of vectors dot product required for per-token quantization.
Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM). We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ because a $K_{j}$ only corresponds to one warp. Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$. According to the layout requirement of this instruction, we find that the $Q_{w}[8\times(n\%8)]$ could share one quantization scale, and $K_{j}[8\times(n\%8)]$ and $K_{j}[8\times(n\%8+1)]$ could share one quantization scale, where $n$ is the token index. Such a quantization method is more fine-grained and with no additional overhead. This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value. The per-thread quantization process can be formulated as follows.
$i_{\delta q}=$ $\lfloor(n*8*c_{w}/b_{q})\rfloor$
$q_{i}[i_{\delta q}]=$$\{8\times(n\%8)+\lfloor(n*c_{w}/b_{q})\rfloor*b_{q}/c_{w}\},~{}~{}n\in[0,N]$
$\delta_{Q}[i_{\delta q}]=$ $\frac{\max(\left|\,Q[q_{i}[i_{\delta q}]]\,\right|)}{7}$
$\hat{Q}[q_{i}[i_{\delta q}]]=$ $\left\lceil\frac{Q[q_{i}[i_{\delta q}]]}{\delta_{Q}[i_{\delta q}]}\right\rfloor$ (5)
$i_{\delta k}=$ $\lfloor(n*4/b_{k})\rfloor$
$k_{n}[i_{\delta k}]=$$\{8\times(n\%8)+\lfloor n/b_{k}\rfloor*b_{k}\}\cup\{8\times(n\%8)+1+\lfloor n/b_{k}\rfloor*b_{k}\},~{}~{}n\in[0,N]$
$\delta_{K}[i_{\delta k}]=$ $\frac{\max(\left|\,K[k_{n}[i_{\delta k}]]\,\right|)}{7}$
$\hat{K}[k_{n}[i_{\delta k}]]=$ $\left\lceil\frac{K[k_{n}[i_{\delta k}]]}{\delta_{K}[i_{\delta k}]}\right\rfloor$ (6)
We show an example of per-thread quantization in Figure 6.
Algorithm 1 Implementation of SageAttention2.
  Input: Matrices $Q(\text{FP16}),K(\text{FP16}),V(\text{FP16})\in\mathbb{R}^{N\times d}$, block size $b_{q},b_{kv}$, warp count $c_{w}$.
  Preprocessing: $K=K-\mathrm{mean}(K)$, $\overrightarrow{Q}_{m}=\mathrm{mean}(Q)$, $Q=Q-\overrightarrow{Q}_{m}$, $\Delta S=\text{GEMV}(\overrightarrow{Q}_{m},K^{\top})$.
  Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}}}(Q)$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K)$. //per-thread. ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$. //per-channel.
  Divide $\hat{Q}$ to $T_{m}={N}/{b_{q}}$ blocks $\{\hat{Q}_{i}\}$; divide $\hat{K}$, $\hat{V}$, and $\Delta S$ to $T_{n}={N}/{b_{kv}}$ blocks $\{\hat{K}_{i}\}$, $\{\hat{V}_{i}\}$, and $\{\Delta S_{i}\}$;
  for $i=1$ to $T_{m}$ do
     Load $\hat{Q}_{i}$ and $\delta_{Q}=\delta_{Q}[c_{w}*8i:c_{w}*8(i+1)]$ into an SM ;
     for j in [1, $T_{n}$do
        Load $\hat{K}_{j}$, $\hat{V}_{j}$, and $\delta_{K}=\delta_{K}[4j:4(j+1)]$ into the SM ;
        $w=\text{range}(c_{w}),st=w*c_{w}$
        $S_{i}^{j}[st:st+c_{w}]=\psi^{-1}_{\delta_{Q}\delta_{K}}(\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top}))+\Delta S_{j}$; // Paralleled by $c_{w}$ warps. The dequantization $\psi^{-1}_{\delta_{Q}\delta_{K}}$ is illustrated in Figure 6.
        $m_{i}^{j}=\mathrm{max}(m_{i}^{j-1},\mathrm{rowmax}(S_{i}^{j}))$, ${}\widetilde{P}_{i}^{j}=\mathrm{exp}(S_{i}^{j}-m_{i}^{j})$, ${}l_{i}^{j}=e^{m_{i}^{j-1}-m_{i}^{j}}+\mathrm{rowsum}(\widetilde{P}_{i}^{j})$ ;
        $O_{i}^{j}$(FP22) = $\mathrm{Matmul}((\widetilde{P}_{i}^{j}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
        $O_{i}^{j}$(FP32) = $\mathrm{diag}(e^{m_{i}^{j-1}-m_{i}^{j}})^{-1}O_{i}^{j-1}(FP32)+O_{i}^{j}$(FP22) ;
     end for
     Load $\delta_{V}$ into an SM;
[…]
     Write $O_{i}$ ;
  end for
  $O=\{O_{i}\}$
  return $O$
Table 4: Error of the FP8 Matmul instruction of mma(f8f8f32).
Precision of Accumulated Value E8M13 E8M23
Error compared to FP32 0 ||||mma(f32.f16.f16.f32) - mma(f32.f8.f8.f32)||||
3.4 FP32 MMA Buffer for FP22 Accumulator
To avoid the drawback of SageAttention where FP16 Matmul with an FP16 accumulator for $PV$ is only effective on GPUs such as the RTX4090, we adopt the approach of quantizing $P$ and $V$ to FP8 to leverage the universal acceleration of FP8 Tensor cores. However, we discover that the accumulator for the mma(f32f8f8f32) instruction on the Ada architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are tensors in FP8 data type and $C,D$ are tensors with FP32 data type, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator. As shown in Table 4, when $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the mma(f16f16f32) instruction. However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32). Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
To mitigate this accuracy loss as much as possible, we propose to use an FP32 buffer to accumulate the values of $\widetilde{P}_{i}^{j}V_{j}$ in FP22. Specifically, as shown in Algorithm 1, after each Matmul between the tiled $\widetilde{P}$ and $V$, we store the Matmul result in the FP32 buffer. This approach prevents the FP22 accumulator from accumulating errors across the entire sequence, confining errors to the tiled range instead.
Table 5: Average accuracy using different data types of $(\widetilde{P},V)$ across all layers of a CogvideoX model, where $(Q,K)$ are smoothed.
$Q,K$ $\widetilde{P},V$ Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
INT4 INT8 77.05% 0.5618 0.5044
E5M2 99.20% 0.0905 0.0903
E4M3 99.44% 0.0683 0.0347
FP16 99.45% 0.0649 0.0335
Table 6: Worst accuracy using different data types of $(\widetilde{P},V)$ across all layers of a CogvideoX model, where $(Q,K)$ are smoothed.
$Q,K$ $\widetilde{P},V$ Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
INT4 INT8 19.52% 0.9579 1.4483
E5M2 94.94% 0.2327 0.2361
E4M3 96.70% 0.1956 0.0779
FP16 96.76% 0.1916 0.0775
Table 7: Two kernel implementations of SageAttention2.
      Kernel       $\psi_{Q}(Q)$, $\psi_{K}(K)$       $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
      SageAttn2-4b       INT4, per-thread       FP8, per-block;  FP8, per-channel
      SageAttn2-8b       INT8, per-thread       FP8, per-block;  FP8, per-channel
3.5 Quantization for Q, K, P, V
Quantization for $Q,K$. We propose $\psi_{Q}(Q)$ and $\psi_{K}(K)$ in granularity of per-thread. This is first because per-channel quantization is not feasible since the scale factors of the inner axis of $QK^{\top}$ cannot used to do dequantization (Xiao et al., 2023a). Second, as shown in Table 3 and Table 3, we compare the average and worst-case accuracies of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX. Results indicate that the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than per-block and per-tensor. Furthermore, per-thread quantization incurs less dequantization overhead than per-token, as discussed in Section 3.3.
Quantization for $\widetilde{P},V$. We choose FP8, specifically the E4M3 data type, for $\psi_{P}(\widetilde{P})$ and $\psi_{V}(V)$ for two reasons. First, most GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16. Second, Table 6 and Table 6 show the average and worst accuracy of different data types used for $\widetilde{P},V$ using real $Q,K,V$ across all layers of CogvideoX. We can see that the accuracy of using E4M3 is very close to that of using FP16 and superior to that of E5M2 and INT8. We propose to use $\psi_{P}(\widetilde{P})$ in per-block and $\psi_{V}(V)$ in per-channel for three reasons. First, per-channel quantization for $\widetilde{P}$ and per-token quantization for $V$ are not viable because dequantization requires scale factors of outer axis. Second, $\widetilde{P}=\mathrm{exp}(S_{i}-\mathrm{rowmax}(S_{i}))$, where $S_{i}$ is the Matmul result of a block of $Q$ and $K^{\top}$, the max value in each row of $\widetilde{P}$ is 1. Hence, we can assign a single static scale $s=\frac{1}{448}$ to a block $\widetilde{P}$. Third, per-channel quantization can address the channel-wised outlier of $V$.
Accuracy metrics. We use three metrics to assess the accuracy of quantized attention output $O^{\prime}$ compared to attention output in full-precision $O$: First, we flatten $O^{\prime}$ and $O$ into vectors in the shape of $1\times n$. Then, Cosine similarity: $CosSim=\sum OO^{\prime}/\smash{\sqrt{\sum O^{2}}}\smash{\sqrt{\sum O^{\prime 2}}}$, Relative L1 distance: $L1=\sum|O-O^{\prime}|/\sum|O|$, Root mean square error: $RMSE=\sqrt{(1/n)\sum(O-O^{\prime})^{2}}$.
Figure 7: An example of dot product precison a row of $\widetilde{P}$ and a column of $V$ presented by FP22 data type.
3.6 An Optional Technique: Smooth V
We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases:
$\gamma(V)=V-\mathrm{mean}(V),$ $~{}~{}~{}~{}\mathrm{mean}(V)=\overrightarrow{V_{m}}$
$\phi_{V}(V)=$ $\psi_{V}\circ\gamma,~{}\overrightarrow{V_{m}}$ (7)
As shown in Figure 7, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9. Consequently, the values of $\widetilde{P}V$ could be quite large. However, the floating-point number representation range is not uniform—it is denser near zero. Therefore, by subtracting the mean along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision. We believe such an explanation and strategy could address lots of issues [^1] https://github.com/triton-lang/triton/issues/4476, https://github.com/triton-lang/triton/issues/5065 reported by the community. Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$. In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$. For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero. This makes the representation of FP22 more accurate. Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
Remark. This technique is optional and not employed in our main experiments. This is because its benefits are marginal, only significantly enhancing attention accuracy primarily when $V$ exhibits channel-wise outliers, which are absent in some models, such as $V$ in Llama3.1 (see Figure 4).
Figure 8: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=64).
Figure 9: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=128).
Figure 10: Speed comparison between SageAttention2 and baselines (L20, headdim=64).
Figure 11: Speed comparison between SageAttention2 and baselines (L20, headdim=128).
4 Experiment
4.1 Setup
Models. We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation. Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2024), HunyuanVideo (Weijie Kong & Jie Jiang, 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
Datasets. Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2023) for evaluating bilingual, multitask, and comprehensive assessment of long context understanding capabilities. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets. Text-to-image models are assessed on MJHQ-30K (Li et al., 2024). TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
Metrics. For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2023). For text-to-video models, following Zhao et al. (2024), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference. For TIMM, we use classification accuracy.
Implemetation details. We implement our attention kernels using CUDA and conduct experiments on Ubuntu 22.04 servers with RTX4090 and L20 GPUs.
Baselines. (1) SmoothAttn. Following Qserve (Lin et al., 2024), we apply smooth quant for matrixes $Q,K$ with smoothing factor $\alpha=0.5$. (2) HadmdAttn. Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for matrixes $Q,K$ before INT4 quantization. (3) SageAttention (Zhang et al., 2024) uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$.
Figure 12: Comparison examples from HunyuanVideo, prompts are sampled from open-sora prompt sets.
Figure 13: Comparison examples from CogvideoX (2B), prompts are sampled from open-sora prompt sets.
Table 8: Average accuracy across all layers of CogvideoX using different smooth methods.
Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
None 80.04% 0.3906 0.2223
HadmdAttn 79.77% 0.3782 0.2180
SmoothAttn 90.21% 0.3383 0.1952
Smooth K (Ours) 98.07% 0.1493 0.0743
Smooth Q (Ours) 98.30% 0.1250 0.0712
SageAttn2-4b 99.46% 0.0648 0.0334
Table 9: Worst accuracy across all layers of CogvideoX using different smooth methods.
Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
None 4.83% 0.9979 0.7784
HadmdAttn 4.85% 0.9978 0.7785
SmoothAttn 64.49% 0.9262 0.7204
Smooth K (Ours) 90.86% 0.3565 0.1464
Smooth Q (Ours) 93.10% 0.2989 0.2195
SageAttn2-4b 96.71% 0.1956 0.0779
Table 10: End-to-end metrics across text and video generation models. - means the model does not support this dataset evaluation. ✗ indicates an inability to generate results for evaluation.
Model Attention WikiText (Ppl.) $\downarrow$ Lambda (Acc.) $\uparrow$ MMLU (Acc.) $\uparrow$ Longbench $\uparrow$
Llama2 Full-Precision 5.823 0.886 0.439 -
HadmdAttn 6.771 0.860 0.360 -
SmoothAttn 6.717 0.867 0.392 -
SageAttn2-4b 5.912 0.881 0.428 -
SageAttn2-8b 5.828 0.886 0.438 -
SageAttention 5.824 0.887 0.439 -
Llama3.1 Full-Precision 6.013 0.815 0.635 49.40
HadmdAttn 7.872 0.762 0.500 44.07
SmoothAttn 7.180 0.783 0.541 44.69
SageAttn2-4b 6.256 0.798 0.607 48.79
SageAttn2-8b 6.019 0.811 0.634 49.59
SageAttention 6.017 0.812 0.634 49.55
GLM4 Full-Precision 7.241 0.432 0.743 49.78
HadmdAttn 7.989 0.435 0.669 45.97
SmoothAttn 8.943 0.449 0.592 42.20
SageAttn2-4b 7.352 0.433 0.725 49.23
SageAttn2-8b 7.242 0.432 0.745 49.60
SageAttention 7.243 0.433 0.744 49.79
Model Attention CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$
CogvideoX (2B) Full-Precision 0.1836 0.9975 77.605 75.360 3.006
HadmdAttn 0.1742 0.9877 29.780 23.985 0.499
SmoothAttn 0.1741 0.9870 41.703 47.043 0.624
SageAttn2-4b 0.1821 0.9973 77.368 74.906 2.603
SageAttn2-8b 0.1829 0.9977 76.532 74.281 2.941
SageAttention 0.1833 0.9976 76.997 71.360 2.988
CogvideoX (1.5-5B) Full-Precision 0.1778 0.9979 70.231 70.928 2.507
HadmdAttn 0.1576 0.9933 8.990 2.299
SmoothAttn 0.1559 0.9950 8.812 2.277
SageAttn2-4b 0.1721 0.9978 57.729 52.989 2.884
SageAttn2-8b 0.1775 0.9980 69.492 74.415 2.487
SageAttention
HunyuanVideo Full-Precision 0.1783 0.9995 82.516 75.934 0.604
HadmdAttn 0.1727 0.9989 7.514 0.762 0.175
SmoothAttn 0.1739 0.9988 6.987 0.609 0.148
SageAttn2-4b 0.1751 0.9995 81.478 65.371 0.610
SageAttn2-8b 0.1782 0.9996 81.786 75.354 0.586
SageAttention 0.1786 0.9995 82.496 79.843 0.597
Mochi Full-Precision 0.1798 0.9986 45.549 65.416 1.266
HadmdAttn 0.1733 0.9980 9.053 25.133 0.704
SmoothAttn 0.1687 0.9978 3.383 3.480 0.241
SageAttn2-4b 0.1783 0.9986 35.955 43.735 1.137
SageAttn2-8b 0.1797 0.9986 46.760 64.901 1.255
SageAttention 0.1800 0.9987 48.707 63.763 1.269
Table 11: Table 10 continued. End-to-end metrics across image generation and classification models.
        Model         Attention        FID $\downarrow$        sFID $\downarrow$        CLIP $\uparrow$        IR $\uparrow$
        Flux         Full-Precision        10.960        16.648        26.180        1.009
        HadmdAttn        11.353        18.495        26.123        0.965
        SmoothAttn        11.149        19.017        26.109        0.959
        SageAttn2-4b        10.577        17.497        26.141        0.998
        SageAttn2-8b        10.927        16.723        26.175        1.009
        SageAttention        10.944        16.641        26.171        1.008
               Stable-Dif        fusion3.5         Full-Precision        14.105        15.646        25.505        0.902
        HadmdAttn        14.259        15.909        25.513        0.886
        SmoothAttn        14.161        15.649        25.510        0.887
        SageAttn2-4b        14.097        15.397        25.487        0.895
        SageAttn2-8b        14.106        15.647        25.499        0.901
        SageAttention        14.140        15.678        25.503        0.902
Model Attention ImageNet (Acc.) $\uparrow$ Sketch (Acc.) $\uparrow$ ImageNet-r (Acc.) $\uparrow$
TIMM Full-Precision 84.79% 45.32% 59.55%
HadmdAttn 84.50% 44.89% 58.80%
SmoothAttn 84.40% 44.68% 58.73%
SageAttn2-4b 86.67% 45.24% 59.29%
SageAttn2-8b 84.79% 45.39% 59.57%
SageAttention 84.74% 45.38% 59.95%
4.2 Speed and Accuracy of Kernels
Speed. We conduct experiments to compare the Speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017). Specifically, Figure 8 and Figure 9, show the speed of SageAttention2 and baselines across varying sequence lengths on RTX4090. These results indicate that SageAttn2-4b and SageAttn2-8b are approximately 3x and 2.7x faster than FlashAttention2, and about 4.5x and 4x faster than xformers, respectively.
Accuracy. Table 9 and Table 9 show that the average and worst accuracy of different methods with INT4 $Q,K$ and FP8 $P,V$ across all layers of CogvideoX. The results indicate that the accuracy of SageAttn2-4b is superior to other baselines.
Table 12: End-to-end generation latency using SageAttention2 (The latency of Llama3.1 is the time to first token generation using different sequence lengths).
Model GPU Original SageAttn2-8b SageAttn2-4b
CogvideoX (2B) RTX4090 86 s 54 s 52 s
CogvideoX (1.5-5B) RTX4090 1040 s 577 s 555 s
HunyuanVideo L20 2221 s 1486 s 1435 s
Mochi L20 2336 s 1316 s 1190 s
Llama3.1 (48K tokens) RTX4090 9.2 s 5.7 s 5.6 s
Llama3.1 (100K tokens) L20 39.9 s 25.4 s 23.2 s
4.3 End-to-end Performance
Metrics loss. We assessed the end-to-end metrics of various models using SageAttention2 compared to using attention in full precision. Detailed evaluation results are presented in Table 10 for Llama2, Llama3.1, GLM4, CogvideoX (2B), CogvideoX (1.5-5B), HunyuanVideo, Mochi, Flux, Stable-Diffusion3.5, and TIMM, respectively. The results indicate that SageAttn2-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models. Additionally, SageAttn2-8b incurs almost no metrics loss across various models.
Visible examples. Figure12 and Figure 13 show some visible comparision examples from HunyuanVideo and CogvideoX (1.5-5B). We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences compared to full-precision attention but is much better than the baselines.
End-to-end speedup. We compared the original generation latency and the latency using SageAttention2 for models with long sequence lengths in Table 12, observing significant speedup effects. For instance, SageAttention2 achieved a 1.8x speedup in CogvideoX (1.5-5B) without any metrics loss (SageAttn2-8b). SageAttn2-4b further accelerated these models but with a little metrics loss.
Table 13: An accuracy example on real tensors of CogvideoX model with or without smoothing $V$.
Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
98.25% 0.1980 0.2387
99.75% 0.0406 0.0773
4.4 Ablation Study
Overhead of smoothing Q, K. Note that the overhead of smoothing $Q,K$ only contains $Q-\overrightarrow{Q_{m}}$, $K-\mathrm{mean}(K)$, and $\overrightarrow{Q_{m}}K^{\top}$. The speed overhead of smoothing $Q,K$ accounts for about 3.5% of the attention kernel.
Benefit of smoothing V. Table 13 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$. It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
5 Conclusion and Future Work
We introduce SageAttention2, an efficient and accurate quantization method for attention. First, we propose to quantize matrixes $(Q,K)$ to INT4 in a thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$. Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$. Our implementation is faster than FlashAttention2 and xformers by approximately 3x and 5x on RTX4090, respectively. Extensive testing confirms that our approach maintains end-to-end metrics across various models, including language, image, and video generation.
Future Work. We leave the implementation of FP8 MatMul with an FP16 accumulator for $\widetilde{P}V$ and SageAttention2 on the Hopper architecture for future work.